OpenStudy (anonymous):

Between what 2 consecutive integers is -sqrt52? A) -5 and -6 B) -6 and -7 C) -7 and -8 D) -8 and -9

OpenStudy (studygurl14):

find each of the numbers in the answer choices squared. The answer is the two number whose square 52 is inbetween. For example... \(\large 2^2 = 4\) and \(\large 3^2=9\). Thus, \(\large 2<\sqrt{6}<3\)

OpenStudy (anonymous):

So it would be C @StudyGurl14